# Thompson Sampling-Based Learning and Control for Unknown Dynamic Systems

### v1

@@ IV-D Analysis of Control Regret
he upper bound of this regret function is provided in the following theorem.
@@
$\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}$
[…]
@@
$+\int_{\mathcal{B}^{c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J}).$ (71)
@@
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})).$ (72)
@@
$\int_{\mathcal{B}^{c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$
@@
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+2\bar{M}_{\bar{\mathcal{J}}}P_{0}\exp(-t\epsilon_{L})$
$\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\!+\!\left[2\bar{M}_{\bar{\mathcal{J}}}\!-\!\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L}).$
@@
$=$ $\epsilon_{J}\frac{\left|J^{*}({\bm{g}}_{t})-J^{*}({\bm{g}}^{*})\right|}{|J^{*}({\bm{g}}_{t})||J^{*}({\bm{g}}^{*})|}\leq\frac{\epsilon_{J}L_{J}}{\underline{J_{m}}^{2}}\|{\bm{g}}_{t}-{\bm{g}}^{*}\|_{\mathcal{L}_{2}}.$ (76)
@@
$\left|\bar{J}^{*}({\bm{g}}_{t})-\bar{J}^{*}({\bm{g}}^{*})\right|\leq\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}.$ (77)
@@
$\leq$ $\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}.$ (79)
@@
$R_{e}({\bm{g}}_{t})\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$
$+\left[2\bar{M}_{\bar{\mathcal{J}}}-\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L})$
$+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}},$ (80)
@@
Theorem 3 provides an upper bound for the regret of the proposed learning Algorithm 2. The regret is decomposed into three terms. The first term represents the regret caused by the neighborhood $\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})$ used in Theorem 2. Intuitively, Theorem 2 shows that the probability of $\bar{J}^{t}\in\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})$ converges to zero but does not analyze the characteristics of the learned distribution function within the neighborhood. Therefore, the characterization of the convergence of the learned function $\bar{J}^{t}$ in Theorem 2 inevitably introduces regret related to the size of the selected neighborhood, which is a constant depending only on the parameter $\delta_{\bar{J}}$. The second term represents the regret caused by the difference between the learned controller $\bar{\bm{g}}_{t}$ and the potential optimal function ${\bm{g}}^{*}$, which is evaluated in Theorem 1 and is caused by the error introduced by the parameterization of the function space. Similar to the first term, the second term is also a constant that does not change during the learning process. The third term is the changing regret during the learning process, which is caused by the gap between the learned distribution function $F^{t}_{\rm pdf}(\cdot)$ and the true value $F^{*}_{\rm pdf}(\cdot)$. It reflects the control regret caused by the unknown system dynamics and noise.

### v2

@@ IV-D Analysis of Control Regret
The upper bound of this regret function is provided in the following theorem.
@@
$\leq$ $\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}J_{\rm min}^{2}}$
$+\left[2\bar{M}_{\bar{\mathcal{J}}}-\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L}),~\text{a.s.},$
@@
$+\int_{\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J}).$ (78)
@@
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})).$ (79)
@@
$\int_{\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$
@@
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+2\bar{M}_{\bar{\mathcal{J}}}P_{0}\exp(-t\epsilon_{L})$
$\leq$ $\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\!+\!\left[2\bar{M}_{\bar{\mathcal{J}}}\!-\!\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L}).$
@@
$=$ $\epsilon_{J}\frac{\left|J^{*}({\bm{g}}_{t})-J^{*}({\bm{g}}^{*})\right|}{|J^{*}({\bm{g}}_{t})||J^{*}({\bm{g}}^{*})|}\leq\frac{\epsilon_{J}L_{J}}{J_{\rm min}^{2}}\|{\bm{g}}_{t}-{\bm{g}}^{*}\|_{\mathcal{L}_{2}}.$ (83)
@@
$\left|\bar{J}^{*}({\bm{g}}_{t})-\bar{J}^{*}({\bm{g}}^{*})\right|\leq\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}J_{\rm min}^{2}}.$ (84)
@@
$\leq$ $\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}J_{\rm min}^{2}}.$ (86)
@@
$R_{e}({\bm{g}}_{t})\leq$ $\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$
$+\left[2\bar{M}_{\bar{\mathcal{J}}}-\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L})$
$+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}J_{\rm min}^{2}},$ (87)
@@
In Theorem 3, the regret bound contains the term $\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$ and the stationary component $\big[2\bar{M}_{\bar{\mathcal{J}}}-\bar{J}_{\rm max}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\big]P_{0}\exp(-t\epsilon_{L})$. An increase in $\delta_{\bar{J}}$ increases the neighborhood measure $v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$ and also increases $\epsilon_{L}$ (see the discussions in Section IV.C). The overall effect on the regret bound is therefore a tradeoff between a larger stationary component and a faster-decaying transient, which depends on the system considered and the time horizon $t$ of the TS-based ALC.